OpenStudy (anonymous):

-3x+6y=18 in slope intercept form

OpenStudy (anonymous):

what are you trying to do?

OpenStudy (anonymous):

Put it in y=mx+b form. The original problem was -3x+2y=6

OpenStudy (anonymous):

well if you have 2y=3x+6 you would have to divide both sides by 2: 2y/2=(3x+6)/2 and this would give you y=1.5x+3

OpenStudy (anonymous):

make sure you don't just divide the 6 by 2 you have to divide the entire right side of the equation by two :)

OpenStudy (anonymous):

Scratch that, wrong problem

OpenStudy (anonymous):

alright haha sorry :P

OpenStudy (anonymous):

It's okay I did what you said for the other problem and got it right